OpenStudy (anonymous):

Simplify: (5.66 x 103)(3.3 x 108) 1.8678 x 1012 1.8678 x 1011 1.8678 x 1024 1.8678 x 1025

OpenStudy (unklerhaukus):

\[(5.66 \times 10^3)(3.3 \times 10^8) \]\[=5.66\times3.3\times10^3\times10^8\]\[=\left(5.66\times3+5.66\times0.3\right)\times10^{3+8}\]\[=(16.98+1.698)\times10^{3+8}\] \[=\cdots\]

OpenStudy (anonymous):

i got the second one am i right

OpenStudy (unklerhaukus):

yes